# Reexamining the strange metal charge response with transmission inelastic electron scattering

### v1

@@ I Introduction
The primary challenge for HTSC, and a broad class of other strongly correlated materials, is therefore to understand the strange metal itself. The inability to account for this state, whose significance extends far beyond superconductivity, has been described as an intellectual crisis [64].
@@
Figure 1: (a) Loss function of Bi-2212 at various momenta, measured with momentum-resolved T-EELS as reported in Ref. [41] (blue line), Ref. [61] (orange line), Ref. [40] (purple line), compared to our current study (black line). (b) Low-energy loss function of Bi-2212 measured with momentum-integrated T-EELS from Ref. [54] (blue line) compared to our current study (black line).(c) Same comparison as (b) but using the data from Ref. [53]. (d) Loss function of Bi-2212 at $q=0$ extracted from several representative IR optics studies [10, 24, 49, 45, 32].
@@ II Previous studies of the charge response of Bi-2212
Focusing on the low-energy portion of the spectrum (0–2 eV), infrared optics studies of Bi-2212 at $q\approx 0$ likewise reported a plasmon excitation that closely resembles the feature observed in T-EELS at the smallest measured momentum [10, 24], albeit at a somewhat different energy. The loss functions extracted from these infrared measurements are shown in Fig. 1(d), labeled “Bozovic 1990” and “Humlicek 1990.”
@@
It is important to note that Ref. [32] reported an unexpectedly weak doping dependence of the 1 eV plasmon in Bi-2212 across samples with superconducting transition temperatures ranging from $T_{c}=60$ K on the underdoped side to $T_{c}=58$ K on the overdoped side. Over this range, the hole concentration in Bi-2212 varies from $p=0.11$ to p=0.24p=0.24—more than a factor of two [18]. In a conventional metal, such a change in carrier density would be expected to result in a plasma frequency shift of more than 40% [3]. Instead, the observed shift was less than 5%, making it difficult to understand the 1 eV excitation as a plasmon in the usual sense. Moreover, this result rules out variations in carrier density as the origin of the differing plasma frequencies shown in Fig. 1(d), suggesting that those discrepancies instead arise from other forms of experimental uncertainty.
@@
A subsequent experiment with improved energy and momentum resolution, $\Delta E=$ 25 meV and $\Delta q\approx 0.11$ Å-1, again failed to detect a 1 eV plasmon in Bi-2212 [54]. Both Tohoku measurements are shown in Fig. 1(b), labeled “Terauchi 1995” and “Terauchi 1999,” respectively.
Most recently, the charge response of Bi-2212 has been investigated using reflection momentum-resolved EELS (M-EELS) [36]. These measurements achieve an energy resolution of 4 meV and a momentum resolution of 0.02 Å-1 [36, 25, 14]. The results are quantitatively consistent with earlier reflection HR-EELS measurements, which claimed an energy resolution of 30 meV and momentum resolution of 0.01 Å-1 [48]. Moreover, in the $q\to 0$ limit, the M-EELS data were shown to agree quantitatively with infrared optics measurements [14, 32].
@@
These discrepancies in the measured charge response of Bi-2212 are deeply concerning. The strange metal remains one of the great unsolved problems for 21st century science. Understanding its origin, and whether its phenomenology adheres to a marginal Fermi liquid picture or something else, ultimately depends on an accurate knowledge of the dynamic charge susceptibility, $\chi^{\prime\prime}(q,\omega)$, at finite momentum, $q$. The inconsistency of experimental results therefore poses a serious problem. Some superficial difference between reflection M-EELS and transmission EELS data is expected because M-EELS probes a surface response whose connection to the bulk susceptibility has only recently been clarified [14]. However, mutually inconsistent T-EELS measurements on the same material represent the core impediment to progress on this critical problem.
@@
A third possibility is differences in momentum resolution. While the studies in Refs. [53, 54] had substantially better energy resolution than the early T-EELS experiments [41, 61, 40], their momentum resolutions of $\Delta q\sim 0.3$ Å-1 and $\Delta q\sim 0.11$ Å-1, respectively, were significantly worse. As a result, these measurements effectively integrated over a significant portion of the plasmon dispersion curve, which would broaden and possibly obscure the excitation. Nevertheless, the plasmon peak in Ref. [41] is clearly visible all the way up to $q=0.35$ Å-1. Therefore, even with this degree of momentum averaging, a well-defined plasmon peak should still have been observable in the data [26].
@@ III Aim of this work
In light of the above inconsistencies, it is necessary to revisit transmission EELS measurements of the finite-$q$ charge response of strange metals, beginning with Bi-2212, with a focus on achieving simultaneous high energy and momentum resolution while minimizing data processing and corrective procedures. Over the past several decades, EELS instrumentation has undergone transformational advances, making it possible to attain meV-scale energy resolution together with high momentum resolution in collimated-beam STEM mode [29, 50, 15]. In parallel, sample preparation techniques for transmission electron microscopy have improved dramatically to meet the needs of aberration-corrected, high-resolution scanning TEM (STEM) studies with sub-Å spatial resolution. Developments driven by two-dimensional materials research have further enabled exceptional control over crystalline flakes down to monolayer thickness, along with highly reproducible preparation protocols. At the same time, advances in detector efficiency now allow measurements at substantially lower electron doses, reducing beam-induced damage while improving sensitivity to weak signals. Finally, the quality and homogeneity of cuprate single crystals have improved significantly over the past two decades [20].
@@ V Results
Figure 4(b) focuses on the low-energy region (0-1.7 eV), corresponding to the shaded area in Fig. 4(a). Data from all ten measurements listed in Table 1 are shown to assess the reproducibility. At the lowest momentum, $q=0.01$ Å-1, the spectra are dominated by the quasielastic peak due to proximity to the bright-field disk, which obscures inelastic features. At $q=0.05$ Å-1, a plasmon-like feature becomes visible, although the degree of interference from the elastic line varies strongly with the momentum resolution. In the measurements with the worst momentum resolution (i.e. D and H), there is no clear maximum. In the measurements with the best momentum resolution, a clear maximum is observed, consistent with a plasmon; however, its linewidth is comparable to its energy, indicating that it is not a well-defined excitation. This feature persists at $q=0.1$ Å-1 and $q=0.15$ Å-1, but shows no clear dispersion, in contrast to expectations for a Fermi-liquid plasmon (see, e.g., Fig. 3(a)). For momenta $q>0.15$ Å-1, the peak is no longer discernible and the spectra evolve into a broad continuum. This behavior differs markedly from Ref. [41], which reported a dispersing plasmon persisting up to momenta $q=0.45$ Å-1 or higher.
@@
In Figs. 1(b)–(c), we compare our results with the high-energy-resolution, momentum-integrated studies of “Terauchi 1999” [54] and “Terauchi 1995” [53], which achieved energy resolutions of 25 meV and 39 meV, respectively. Although these studies achieved substantially higher energy resolution than earlier work, their momentum resolution was comparatively poor—0.11 Å-1 in the 1999 study and 0.3 Å-1 in the 1995 study. To enable a direct comparison, we artificially broaden our momentum resolution by summing spectra over momenta up to 0.11 Å-1 and 0.3 Å-1, respectively, weighting the data to account for the circular selection aperture used in those studies. The resulting spectra are shown in Figs. 1(b) and (c).
@@
The agreement in the low-energy range in Figs. 1(b)-(c) arises, in part, because the quasielastic peak was retained in both of these early studies. Notably, Terauchi (1995) [53] demonstrated that subtracting the elastic line from these spectra produced a plasmon-like feature that was not present in the raw data. A key takeaway from these studies is that elastic-line subtraction can give the appearance of a plasmon even when none exists in the measured spectra.
@@ VI Conclusions
We report the first transmission EELS study of the charge response of Bi-2212 that combines high energy resolution ($\Delta E\approx$ 39 meV) with high momentum resolution ($\Delta q\approx 0.014$Å-1). To address historical reproducibility issues in this field, we repeated the experiment ten times on five different Bi-2212 flakes (Table 1), validated the setup using a standard aluminum reference (Fig. 3(a)), and made a quantitative comparison with previous studies dating back to 1989. A broad, plasmon-like feature is observed at $q<0.15$ Å-1; however, we do not observe the RPA-like dispersion reported by Nücker et al. (1989) [41]. Instead, the plasmon simply damps out at these momenta. The pronounced plasmon reported at large $q$ in that early study may be an artifact of subtraction of the zero-loss line, as previously noted by the Tohoku group [53].
@@
One limitation of the present T-EELS experiments is that, unlike M-EELS [14], we were unable to access momenta small enough to reach the optical limit ($q<0.05$ Å-1), where the spectra are obscured by the tails of the bright-field disk. Further work is therefore needed to assess the quantitative consistency of T-EELS measurements with infrared-optics studies, such as those shown in Fig. 1(d).
@@ VII Acknowledgments
We gratefully acknowledge M. Terauchi, J. Zaanen, E. Huang, B. Bradlyn, and P.W. Phillips for helpful discussions, and D. van der Marel for supplying infrared data. This work was supported by the Center for Quantum Sensing and Quantum Materials, an Energy Frontier Research Center funded by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award no. DE-SC0021238. Momentum resolved transmission EELS measurements at the Center for Nanophase Materials Sciences (CNMS) at Oak Ridge National Laboratory were supported by UT-Battelle, LLC under DOE BES contract no. DE-AC05-00OR22725, as well as by the Laboratory Directed Research and Development Program of ORNL. P.A. acknowledges additional support from the EPiQS program of the Gordon and Betty Moore Foundation under Grant No. GBMF9452.

### v2

@@ I Introduction
The primary challenge for HTSC, and a broad class of other strongly correlated materials, is therefore to understand the strange metal itself. The inability to account for this state, whose relevance extends far beyond superconductivity, has been described as an intellectual crisis [64].
@@
Figure 1: (a) Loss function of Bi-2212 at various momenta, measured with momentum-resolved T-EELS as reported in Ref. [41] (blue line), Ref. [61] (orange line), Ref. [40] (purple line), compared to our current study (black line). (b) Low-energy loss function of Bi-2212 measured with momentum-integrated T-EELS from Refs. [54] (blue line) and  [53] (orange line) compared to our current study (black line). (c) Loss function of Bi-2212 at $q=0$ extracted from several representative IR optics studies [10, 24, 49, 45, 32].
@@ II Previous studies of the charge response of Bi-2212
Focusing on the low-energy portion of the spectrum (0–2 eV), infrared optics studies of Bi-2212 at $q\approx 0$ likewise reported a plasmon excitation that closely resembles the feature observed in T-EELS at the smallest measured momentum [10, 24], albeit at a somewhat different energy. The loss functions extracted from these infrared measurements are shown in Fig. 1(c), labeled “Bozovic 1990” and “Humlicek 1990.”
@@
It is important to note that Ref. [32] reported an unexpectedly weak doping dependence of the 1 eV plasmon in Bi-2212 across samples with superconducting transition temperatures ranging from $T_{c}=60$ K on the underdoped side to $T_{c}=58$ K on the overdoped side. Over this range, the hole concentration in Bi-2212 varies from $p=0.11$ to p=0.24p=0.24—more than a factor of two [18]. In a conventional metal, such a change in carrier density would be expected to result in a plasma frequency shift of more than 40% [3]. Instead, the observed shift was less than 5%, making it difficult to understand the 1 eV excitation as a plasmon in the usual sense. Moreover, this result rules out variations in carrier density as the origin of the differing plasma frequencies shown in Fig. 1(c), suggesting that those discrepancies instead arise from other forms of experimental uncertainty.
@@
A subsequent experiment with improved energy resolution, $\Delta E=$ 25 meV and $\Delta q\approx 0.3$ Å-1, again failed to detect a 1 eV plasmon in Bi-2212 [54]. Both Tohoku measurements are shown in Fig. 1(b), labeled “Terauchi 1995” and “Terauchi 1999,” respectively.
Most recently, the charge response of Bi-2212 has been investigated using reflection momentum-resolved EELS (M-EELS) [36]. These measurements achieve an energy resolution of 4 meV and a momentum resolution of 0.02 Å-1 [36, 25, 14]. The results are quantitatively consistent with earlier reflection HR-EELS measurements, which stated an energy resolution of 30 meV and momentum resolution of 0.01 Å-1 [48]. Moreover, in the $q\to 0$ limit, the M-EELS data were shown to agree quantitatively with infrared optics measurements [14, 32].
@@
These discrepancies in the measured charge response of Bi-2212 are deeply concerning. The strange metal remains one of the great unsolved problems for 21st century science. Understanding its origin, and whether its phenomenology adheres to a marginal Fermi liquid picture or something else, ultimately depends on an accurate knowledge of the dynamic charge susceptibility, $\chi^{\prime\prime}(q,\omega)$, at finite momentum, $q$. The inconsistency of experimental results therefore poses a serious problem. Some superficial difference between reflection M-EELS and transmission EELS data is expected because M-EELS probes a surface response whose connection to the bulk susceptibility has only recently been quantified [14]. However, mutually inconsistent T-EELS measurements on the same material represent the core impediment to progress on this critical problem.
@@
A third possibility is differences in momentum resolution. While the studies in Refs. [53, 54] had substantially better energy resolution than the early T-EELS experiments [41, 61, 40], their momentum resolution of $\Delta q\sim 0.3$ Å-1 is significantly worse. As a result, these measurements effectively integrated over a significant portion of the plasmon dispersion curve, which could broaden and possibly obscure the excitation. Nevertheless, the plasmon peak in Ref. [41] is clearly visible all the way up to $q=0.35$ Å-1. Therefore, even with this degree of momentum averaging, a well-defined plasmon peak should still have been observable in the data [26].
@@ III Aim of this work
In light of the above inconsistencies, it is necessary to revisit transmission EELS measurements of the finite-$q$ charge response of strange metals, beginning with Bi-2212, with a focus on achieving simultaneous high energy and momentum resolution while minimizing data processing and corrective procedures. Over the past several decades, EELS instrumentation has undergone transformational advances, making it possible to attain meV-scale energy resolution together with high momentum resolution in collimated-beam STEM mode [29, 50, 15]. In parallel, sample preparation techniques for transmission electron microscopy have improved dramatically to meet the needs of aberration-corrected, high-resolution scanning TEM (STEM) studies with sub-Å spatial resolution. Developments driven by two-dimensional materials research have further enabled unprecedented control over crystalline flakes down to monolayer thickness, along with highly reproducible preparation protocols. At the same time, advances in detector efficiency now allow measurements at substantially lower electron doses, reducing beam-induced damage while improving sensitivity to weak signals. Finally, the quality and homogeneity of cuprate single crystals have improved significantly over the past two decades [20].
@@ V Results
Figure 4(b) focuses on the low-energy region (0-1.7 eV), corresponding to the shaded area in Fig. 4(a). Data from all ten measurements listed in Table 1 are shown to assess the reproducibility. At the lowest momentum, $q=0.01$ Å-1, the spectra are dominated by the quasielastic peak due to proximity to the bright-field disk, which obscures inelastic features. At $q=0.05$ Å-1, a plasmon-like feature becomes visible, although the degree of interference from the elastic line varies strongly with the momentum resolution. In the measurements with the worst momentum resolution (i.e. D and H), there is no clear maximum. In the measurements with the best momentum resolution, a clear maximum is observed, consistent with a plasmon; however, its linewidth is comparable to its energy, indicating that it is only a marginally well-defined excitation. This feature persists at $q=0.1$ Å-1 and $q=0.15$ Å-1, but shows no clear dispersion, in contrast to expectations for a Fermi-liquid plasmon (see, e.g., Fig. 3(a)). For momenta $q>0.15$ Å-1, the peak is no longer discernible and the spectra evolve into a broad continuum. This behavior differs markedly from Ref. [41], which reported a dispersing plasmon persisting up to momenta $q=0.45$ Å-1 or higher.
@@
In Figs. 1(b), we compare our results with the high-energy-resolution, momentum-integrated studies of “Terauchi 1999” [54] and “Terauchi 1995” [53], which achieved energy resolutions of 25 meV and 39 meV, respectively. Although these studies achieved substantially higher energy resolution than earlier work, their momentum resolution was comparatively poor—0.3 Å-1 in both the 1995 and 1999 studies. To enable a direct comparison, we artificially broaden our momentum resolution by summing spectra over momenta up to 0.3 Å-1, weighting the data to account for the circular selection aperture used in those studies. The resulting spectrum is shown in Fig. 1(b).
@@
The agreement in the low-energy range in Figs. 1(b) arises, in part, because the quasielastic peak was retained in both of these early studies. Notably, Terauchi (1995) [53] demonstrated that subtracting the elastic line from these spectra produced a plasmon-like feature that was not present in the raw data. A key takeaway from these studies is that elastic-line subtraction can give the appearance of a plasmon even when none exists in the measured spectra.
@@ VI Conclusions
We report the first transmission EELS study of the charge response of Bi-2212 that combines high energy resolution ($\Delta E\approx$ 39 meV) with high momentum resolution ($\Delta q\approx 0.014$Å-1). To address historical reproducibility issues in this field, we repeated the experiment ten times on five different Bi-2212 flakes (Table 1), validated the setup using a standard aluminum reference (Fig. 3(a)), and made a quantitative comparison with previous studies dating back to 1989. A broad, plasmon-like feature is observed at $q<0.15$ Å-1; however, we do not observe the RPA-like dispersion reported by Nücker et al. (1989) [41]. Instead, the plasmon simply damps out at larger momenta. The pronounced plasmon reported at large $q$ in that early study may be an artifact of subtraction of the zero-loss line, as previously noted by the Tohoku group [53].
@@
One limitation of the present T-EELS experiments is that, unlike M-EELS [14], we were unable to access momenta small enough to reach the optical limit ($q<0.05$ Å-1), where the spectra are obscured by the tails of the bright-field disk. Further work is therefore needed to assess the quantitative consistency of T-EELS measurements with infrared-optics studies, such as those shown in Fig. 1(c).
@@ VII Acknowledgments
We gratefully acknowledge M. Terauchi, J. Zaanen, E. Huang, B. Bradlyn, and P.W. Phillips for helpful discussions, and D. van der Marel for supplying infrared data. This work was supported by the Center for Quantum Sensing and Quantum Materials, an Energy Frontier Research Center funded by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award no. DE-SC0021238. Momentum resolved transmission EELS measurements at the Center for Nanophase Materials Sciences (CNMS) at Oak Ridge National Laboratory were supported by UT-Battelle, LLC under DOE BES contract no. DE-AC05-00OR22725, as well as by the Laboratory Directed Research and Development Program of ORNL. Some sample preparation instrumentation was supported by NSF MRSEC award DMR-2309037. P.A. acknowledges additional support from the EPiQS program of the Gordon and Betty Moore Foundation under Grant No. GBMF9452.